OpenStudy (anonymous):

divide and simplify x^7+x^6/x

OpenStudy (anonymous):

\[\frac{ x ^{7}+x ^{6} }{ x }\]

OpenStudy (anonymous):

x^7/x=? x^6/x=?

OpenStudy (anonymous):

is the answer x^6+x^5

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

omg I got something right lol